OpenStudy (anonymous):

The half life of the radioactive material radium-226 is 1590 years. (a) A sample of the radium-226 has a mass of 100mg. Find a formula for the mass of radium-226 that remains after t years. (b) FInd the mass after 1000 years correct to the nearest milligram. (c) when will the mass be reduced to 30 mg? (d) At what rate is the mass decreasing when 30 mg remains?

OpenStudy (anonymous):

\[\large A=A _{0}*2^{-t/h}\] A=final amount A_o=initial amount t=time h=half life wanna plug some things in that we know?

OpenStudy (anonymous):

But isn't the formula for decay \[A(t)= C e ^{kt}\] ?

OpenStudy (anonymous):

So then when i plug things in it would be \[100e ^{kt}\] but i dont know what to do after.

OpenStudy (anonymous):

That does look more familiar. Im probably mixing things up, well go with yours.

OpenStudy (anonymous):

One second while I refresh myself I remember doing this last semester

OpenStudy (anonymous):

Okay thanks!

OpenStudy (anonymous):

ok i remember. we know an obvious solution, since the half life is 1590 years and we start with 100mg, after 1590 years we will have 50mg. so basically that gives us some variables we can plug in then solve for k. once we know k well have the formula and can solve for the amount at any given time 50=100e^(1590k) so solve for k in that equation

OpenStudy (anonymous):

so would it be \[\ln \frac{ 1 }{ 5 }\ /1590\] ?

OpenStudy (anonymous):

i mean 1/2

OpenStudy (anonymous):

1/2=e^1590k ln1/2=1590k (ln1/2)/1590=k that is correct. so now that we know that, we can set our equation up \[A(t)=100e^{((\ln1/2)/1590))t}\] so theres the answer to a

OpenStudy (anonymous):

For b, you just wanna plug in 1000 for t, and solve

OpenStudy (anonymous):

Would the answer be? \[a(t)= 1000e ^{\ln \frac{ 1 }{ 2 }/1590 (1000) t}\]\[= 64.7 mg\]

OpenStudy (anonymous):

and sorry i mean 100e

OpenStudy (anonymous):

yep so to the nearest milligram, 65mg c) plug 30 in for the amount and solve for t

OpenStudy (anonymous):

I know I have to take the logarithm of both sides so one side would be ln 30 but how would i take ln from \[100e ^{\frac{ \ln 1/2 }{ 1590}(t)}\]

OpenStudy (anonymous):

give me a second, doing it on paper

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

ok, what you do is this. 30=100e^((ln1/2)/1590)t divide by 100 3/10=e^((ln1/2)/1590)t take ln of both sides, remember ln is natural log and the base is 'e', so the ln and e cancel out ln 3/10=((ln 1/2)/1590)t divide both sides by ((ln(1/2)/1590), when you divide by a fraction you multiply by the reciprocal (ln 3/10)*((1590/ln(1/2))=t wanna finish?

OpenStudy (anonymous):

so t = (1590)(ln 3/10) / (ln 1/2) or can it be simplified even further?

OpenStudy (anonymous):

I think thats as good as it gets. slap that in a calculator to get your number answer

OpenStudy (anonymous):

it would be around 2762 years

OpenStudy (anonymous):

Correct. I'm not sure about d. Do you have any ideas?

OpenStudy (anonymous):

No, I have no idea. Oh well. You've helped me enough. Thanks for your help!

OpenStudy (anonymous):

You're welcome